x and y are real numbers and x>y
x^2-xy+y^2=15
xy+x+y=13
find x^2+6y

Grade:10

1 Answers

Sher Mohammad IIT Delhi
9 years ago
x^2-xy+y^2+3(xy+x+y)=15+3*13=64
(x+y)^2+3(x+y)=64
let x+y=z
z^2+3z-64=0
x+y, 6.63, -.9.63
xy=13-6.63=.37
use the above two equation to calculate x^2+6y